OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3? @MarrBear

OpenStudy (anonymous):

x + 3y = 10 x + 3y = −10 3x + y = 10 3x + y = −10

OpenStudy (anonymous):

Im pretty sure its D.

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1) this is the equation.. m = slope

OpenStudy (anonymous):

But with this question, im not 100%.

OpenStudy (anonymous):

ok cool try this question Jack and Jill had 2 pails of water. Each trip up the hill, Jack and Jill obtained 4 more buckets of water. Which of the graphs below represents Jack and Jill's number of pails over time?

OpenStudy (anonymous):

Wheres he graphs ;o

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

The first one. c:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Anytime c: